OpenStudy (anonymous):

hi I need to get line integral: Integral( (4x+y)dx + (x+4y)dy ) and the curve is y=x^4 from A(-1,1) To B(1,1) please help

hartnn (hartnn):

hi :) since, y= x^4 you can plug in x^4 instead of 'y' in ( (4x+y)dx + (x+4y)dy ) also, for dy, you need to differentiate y=x^4 to get dy have you tried these things ?

OpenStudy (anonymous):

I have parametrized y and x like t=x so y=t*4 then I plugged it into the function and get incorrect answer the correct answer is -2, could you please write how you did it? because when I try to solve it i think i am making everything fine but getting wrong answer thank you

hartnn (hartnn):

ok, i get +2...maybe you could find any sign mistake in my work... x=t --> dx=dt, y=t^4---->dy=4t^3dt \((4t+x^4)dt +(t+4t^4)(4t^3)dt = (4t+x^4+4t^4+16t^7)dt\) \(\large \int \limits_{-1}^1 (4t+x^4+4t^4+16t^7)dt = 0+ \int \limits_{-1}^1 5t^4 dt\) because 't' and t^7 are odd functions of 'x' and when you integrate an odd function from -a to a , you get a 0 \(\large = 5 (t^5/5) = [t^5]_{-1}^1 = 1^5-(-1)^5=1-(-1)=2\)

hartnn (hartnn):

that should be 4t+t^4 everywhere .....typing error

hartnn (hartnn):

thats how t^4+4t^4 becomes 5t^4

OpenStudy (anonymous):

don't think you need a separate parameter here, since you know y as a function of x you can substitute for y and make sure you take the differential of y w.r.t x

hartnn (hartnn):

^correct, thats what i said in my first post, but using parameter is also not wrong, it'll yield same answer...

OpenStudy (anonymous):

thanks lot

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

sorry, didn't read that :)